Sally brought three chocolate bars and a pack of gum and paid $1.75. Jake brought two chocolate bars and four packs of gum and paid $2.00
Find the cost of the chocolate bar and the cost of a pack of gum

Answer by scott8148(6628)

3c+g=175 ... g=175-3c

2c+4g=200 ... substituting gives 2c+4(175-3c)=200 ... 2c+700-12c=200 ... -10c=-500